OpenStudy (anonymous):

Is a triangle whit sides of length 6ft,21ft,23ft, a right triangle ? please guys explain me in to words ,I understand better

OpenStudy (raden):

would the value of 23^2 = 6^2 + 21^2 ? if yes, it is a right triangle...

OpenStudy (anonymous):

is starting 6ft, 21ft,23ft, then I don't know the difference

OpenStudy (calculusxy):

You have to see if the triangle matches the features. A right triangle will match the Pythagorean Theorem meaning a^2=b^2=c^2. Plug in the lengths to the values. The c side will be the longer side.

OpenStudy (anonymous):

it does no show is only a question whit no graph

OpenStudy (raden):

just checking this calculation : 6^2 + 21^2 = 23^2 (is the left side be equal with the right side?) 36 + 441 = 529 477 = 529 (LHS not equal RHS) so, it is not a right triangle